OpenStudy (anonymous):

Concrete can be purchased by the cubic yard. How much will it cost to pour a slab 18ft by 18ft by 4in for a patio if the concrete costs $41.00 per cubic yard?

OpenStudy (kittiwitti1):

18ft•18ft•4in making it equal: 3ft per yd so 18/3=6yd; 36in per yd so 4/36=1/9yd now we get this equation:\[6\times6\times\frac{1}{9}\text{, all in yards}\]After you get the result for that, multiply by $41 per yard and you will get the cost ☺

OpenStudy (anonymous):

Your amazing! Thank you so much!